OpenStudy (anonymous):

6) Graph the circle with center at (-1, 2) and diameter 12. Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (zpupster):

(h,k) is center and equation for circle is (x-h)^2 + (y-k)^2 = r^2 so h =-1 and k =2 the diameter = 12 r or the radius is d/2 12/2 =? can you plug theses in